OpenStudy (anonymous):

If (theta) = arctan1, what is theta? in terms of pi?

OpenStudy (anonymous):

tan(theta)=tan(arctan1) tan(theta)=1 think of your unit circle tan(theta)=y/x if y/x=1, then y=x. you have an isosceles right triangle, the angles of which are each 45º or pi/4 radians

OpenStudy (anonymous):

that is the reference angle. there are other solutions

OpenStudy (anonymous):

theta=pi/4+k*pi, where k is an integer

OpenStudy (anonymous):

I have a similar problem that I can't figure out. if theta = arctan(-1/squareroot 3) what is theta in terms of pi?